A closed-end fund starts the year with a net asset value of $20. By year-end, NAV equals $20.90. At the beginning of the year, the fund is selling at a 4% premium to NAV. By the end of the year, the fund is selling at a 9% discount to NAV. The fund paid year-end distributions of income and capital gains of $2.30.


a. What is the rate of return to an investor in the fund during the year? (Do not round intermediate calculations. Round your answer to 2 decimal places.)


Rate of return             %


b. What would have been the rate of return to an investor who held the same securities as the fund manager during the year? (Round your answer to 2 decimal places.)


Rate of return             %
Answer #1

a)

P0 = NAV0(1+premium) = 20*1.04 = 20.8

P1 = NAV1(1-discount) = 20.90*0.91 = 19.019

Return = (P1+ Distributions)/P0 -1 =  19.019+2.30/20.8 = 2.50%

b)

Return = (NAV1+ Distributions)/NAV0– 1

=20.90+2.30/20 - 1

=16.00%